# Thermodynamics of the FLRW apparent horizon in Einstein-Gauss-Bonnet gravity

### v1

@@ I Introduction
This idea has been applied in the case of cosmological models, where it was proved that the apparent horizon of a Friedmann-Lemaître-Robertson-Walker (FLRW) can be equipped with an entropy proportional to the area of the horizon, from which the Friedmann equations are derived Cai and Kim (2005). Following this approach, in a previous work Sánchez and Quevedo (2023), we assumed that the FLRW apparent horizon itself is a thermodynamic system that can be handled by using the standard relations of classical thermodynamics. As a result, we proved that, in fact, the apparent horizon can be characterized by a fundamental equation, from which all the properties of the system are derived. In particular, we found that the apparent horizon satisfies a dark-energy-like equation of state and satisfies the corresponding laws of thermodynamics when considered as a system surrounded by the fluids of the standard cosmological model.
To find out if the above result can be generalized to other gravity theories, in this work, we study the FLRW spacetime in the Einstein-Gauss-Bonnet (EGB) gravity theory. We will show that, in fact, the FLRW apparent horizon can be handled as a thermodynamic system. The thermodynamic variables and the response functions turn out to depend explicitly on the Gauss-Bonnet (GB) coupling constant and on the number of dimensions. However, the equation of state of the horizon does not depend on the details of the EGB theory.
This work is organized as follows. In Sec. II, we review the main concepts of the EGB theory and present the field equations of the FLRW cosmological model. In Sec. III, we consider the properties of the apparent horizon in the FLRW spacetime, and show that in the EGB theory, the horizon radius is always larger than in Einstein’s gravity. In Sec. IV, we derive the explicit form of the main thermodynamic variables and explore the conditions that the apparent horizon must satisfy for the laws of thermodynamics to be fulfilled. The consequences of this assumption are presented in Sec. V, where it is shown that the apparent horizon satisfies a dark-energy-like equation of state, implying that the thermodynamic evolution of the horizon is isentropic. We argue that this result can lead to a discretization of the GB coupling constant. Finally, in Sec. VI, we sumarize and discuss our results.
@@ II The FLRW spacetime in EGB gravity
The action of the EGB gravity theory can be written as Cai and Kim (2005)
@@
where $\alpha_{0}$ is the GB coupling constant, $\Lambda$ si the cosmological constant, $S_{m}$ represents the action of the matter, and $\mathcal{G}$ is the so-called GB term given by
@@
In four dimensions, the GB term is purely topological, so it does not contribute to the dynamics. What distinguishes the EGB gravity is that, even though the action involves higher-order curvature terms, the equations of motion derived from it contain only second-order derivatives of the metric. Indeed, by varying the action, one obtains the equations of motion
@@
We will consider the FLRW spacetime, which is determined by the metric
@@
where the index $a=0,1$, $d\Omega_{n-1}^{2}$ is the line element of a unit $(n-1)$ dimensional sphere, $k=0,\pm 1$ is the curvature constant, the coordinates $x^{a}$ are defined as $x^{0}=t$ and $x^{1}=r$, and we introduce the notation $R=a(t)r$. Moreover, the gravitational source of the FLRW spacetime is assumed to be determined by the energy-momentum tensor of a perfect fluid
@@ III The FLRW apparent horizon
An important feature of the FLRW spacetime is the existence of a marginally trapped null surface, known as the apparent horizon, characterized by a vanishing expansion. This surface satisfies the condition Faraoni (2011, 2019)
@@
Note that the energy flux vector and the work density depend solely on the energy-momentum tensor and the metric in the $(t,r)$ plane. However, both quantities are independent of the field equations. Therefore, the thermodynamic quantitites $\Psi_{a}$ and $W$ coincide in EGB gravity and in general relativity. The same holds for the surface gravity, as it is defined within the $(t,r)$ plane and depends exclusively on the metric $h_{ab}$.
@@ III.1 Analysis of the apparent horizon radius
which in the limiting case $\alpha=0$ and $n=3$ reduces to the corresponding value in Einstein’s gravity Sánchez and Quevedo (2023)
@@ IV The laws of thermodynamics for the FLRW apparent horizon
which is the entropy of the FLRW apparent horizon in Einstein’s gravity Sánchez and Quevedo (2023).
In Fig. 3, we illustrate the behavior of the entropy (32) in terms of the energy for different values of the coupling constant. We see that for negative values of $\alpha_{0}$, the entropy can become negative, which we interpret as a non-physical behavior. In general, for positive values of $\alpha_{0}$ the entropy is always larger than in the case of Einstein’s gravity. Figure 4 shows the entropy in terms of the energy for a fixed value of $\alpha_{0}$ and different dimensions. In all the cases, the entropy is larger than in Einstein’s gravity.
@@ IV.1 The first law of thermodynamics
In the case of an apparent horizon, the first law can be derived using the conservation of energy, which states that the change of internal energy is equivalent to the amount of energy that crosses the horizon during its evolution Cai and Kim (2005). Then, the first law can be expressed as
$dE=-TdS+WdV.$ (34)

### v3

@@ I Introduction
This idea has been applied in the case of cosmological models, where it was proved that the apparent horizon of a Friedmann-Lemaître-Robertson-Walker (FLRW) can be equipped with an entropy proportional to the area of the horizon, from which the Friedmann equations are derived Cai and Kim (2005). Following this approach, in a previous work Sánchez and Quevedo (2023), we assumed that the FLRW apparent horizon itself is a thermodynamic system that can be handled by using the standard relations of classical thermodynamics. As a result, we proved that, in fact, the apparent horizon can be characterized by a fundamental equation, from which all the properties of the system are derived. In particular, we found that the apparent horizon satisfies a dark-energy-like equation of state and satisfies the corresponding laws of thermodynamics when considered as a system surrounded by the fluids of the standard cosmological model. Other aspects of the application of thermodynamics in cosmology have been analyzed in Nojiri et al. (2025); Odintsov et al. (2025); Paul (2025).
To find out if the above result can be generalized to other gravity theories, in this work, we study the FLRW spacetime in the Einstein-Gauss-Bonnet (EGB) gravity theory. Several interesting thermodynamic aspects of the FLRW spacetime in the EGB theory have been analyzed in Saavedra and Tello-Ortiz (2025); Paul et al. (2024). In this work, we will show that, in fact, the FLRW apparent horizon can be handled as a thermodynamic system. The thermodynamic variables and the response functions turn out to depend explicitly on the Gauss-Bonnet (GB) coupling constant and on the number of dimensions. Remarkably, the equation of state of the horizon turns out to be independent of the details of the EGB theory.
This work is organized as follows. In Sec. II, we review the main concepts of the EGB theory and present and analyze the field equations, which describe the evolution of the FLRW cosmological model. In Sec. III, we consider the properties of the apparent horizon in the FLRW spacetime, and show that in the EGB theory, the horizon radius is always larger than in Einstein’s gravity. In Sec. IV, we derive the explicit form of the main thermodynamic variables and explore the conditions that the apparent horizon must satisfy for the laws of thermodynamics to be fulfilled. The consequences of this assumption are presented in Sec. V, where it is shown that the apparent horizon satisfies a dark-energy-like equation of state, implying that the thermodynamic evolution of the horizon is isentropic. We argue that this result can lead to a discretization of the GB coupling constant. Finally, in Sec. VI, we summarize and discuss our results.
@@ II The FLRW spacetime in EGB gravity
The action of EGB gravity theory can be written as Cai and Kim (2005)
@@
where $\alpha_{0}$ is the GB coupling constant, $\Lambda$ is the cosmological constant, $S_{m}$ represents the action of the matter, and $\mathcal{G}$ is the so-called GB term given by
@@
In four dimensions, the GB term is purely topological and, therefore, it does not contribute to the dynamics. What distinguishes the EGB gravity is that, even though the action involves higher-order curvature terms, the equations of motion derived from it contain only second-order derivatives of the metric. Indeed, it can be shown that by varying the action, one obtains the equations of motion
@@
In this wirk, we will consider the FLRW cosmological model, which is described by the metric
@@
where the index $a=0,1$, $d\Omega_{n-1}^{2}$ is the line element of a unit (n1)(n-1) dimensional sphere, $k=0,\pm 1$ is the curvature constant, the coordinates $x^{a}$ are defined as $x^{0}=t$ and $x^{1}=r$, and we introduce the notation $R=a(t)r$. Moreover, the gravitational source of the FLRW spacetime is assumed to be determined by the energy-momentum tensor of a perfect fluid
@@ III The FLRW apparent horizon
An important feature of the FLRW spacetime is the existence of a marginally trapped null surface, known as the apparent horizon, characterized by a vanishing expansion. This surface satisfies the condition Faraoni (2011, 2015)
@@
Note that the energy flux vector and the work density depend solely on the energy-momentum tensor and the metric in the $(t,r)$ plane. However, both quantities are independent of the field equations. Therefore, in EGB gravity, the thermodynamic quantities $\Psi_{a}$ and $W$ coincide with those of general relativity. The same holds for the surface gravity, as it is defined within the $(t,r)$ plane and depends exclusively on the metric $h_{ab}$.
@@ III.1 Analysis of the apparent horizon radius
which in the limiting case $\alpha=0$ and $n=3$ reduces to the corresponding value in Einstein gravity Sánchez and Quevedo (2023)
@@ IV The laws of thermodynamics for the FLRW apparent horizon
which is the entropy of the FLRW apparent horizon in Einstein gravity Sánchez and Quevedo (2023).
In Fig. 3, we illustrate the behavior of the entropy (32) in terms of the energy for different values of the coupling constant. We see that for negative values of $\alpha_{0}$, the entropy can become negative, which we interpret as a non-physical behavior. In general, for positive values of $\alpha_{0}$ the entropy is always larger than in the case of Einstein gravity. Figure 4 shows the entropy in terms of the energy for a fixed value of $\alpha_{0}$ and different dimensions. In all the cases, the entropy is larger than in Einstein’s gravity.
@@ IV.1 The first law of thermodynamics
In the case of an apparent horizon, the first law can be derived using the law of conservation of energy, which states that the change of internal energy is equivalent to the amount of energy that crosses the horizon during its evolution Cai and Kim (2005). Then, the first law can be expressed as
dE=TdS+WdV.dE=-TdS+WdV. (34)